Donald purchased a bond which carries a coupon rate of 10%. The bond is a 5 million 3 year bond. The bond was purchased in Jan 1 2019 to yield 12%. The interest is payable annually and the principal is to be repaid on maturity date.

The market interest rate
On Dec 31 2019 is 11.5%.
On dec 31 2020 12.5%.
On dec 31 2021 9%

what is the carrying amount of the investment on Dec. 31 2019?
Answer #1

price of bond = present value of annual payments + present value of principal

= (500000*PVFA 12%, 3 years) + (5000000*PVF 12%,3 years)

= (500000*2.40183)+ (5000000*0.71178) = $4759815

carrying amount of investment on Dec. 31 2019 = price of bond + interest expense - interest payment = 4759815+(4759815*11.5%)-(5000000*10%) = $4807194